Two angles are complimentary so there sum is 90 degrees. If the measure of the larger angles is 15 more than two times the smaller angle, find themeasure of the two angles
Answer by rfer(16322)

x+2x+15=90
3x=90-15
3x=75
x=75/3
x=25
2x+15=65